# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ 5.1. Shift invariant kernels, lattices, and the bit-reversed FFT/IFFT
$\eta_{\alpha}(\delta)=\frac{(2\pi)^{2\alpha}}{(-1)^{\alpha+1}(2\alpha)!}B_{2\alpha}(\delta)$
with ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$, ${\mathcal{D}}=\{1,\dots,d\}$, weights $\{\gamma_{\mathfrak{u}}\}_{{\mathfrak{u}}\subseteq{\mathcal{D}}}$, and $B_{\ell}$ denoting the Bernoulli polynomial of degree $\ell$. The corresponding reproducing kernel Hilbert space (RKHS) $H$ is a weighted periodic unanchored Sobolev space of smoothness ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$ with norm
$\lVert f\rVert_{H}^{2}:=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\frac{1}{(2\pi)^{2\lvert{\mathfrak{u}}\rvert}\gamma_{\mathfrak{u}}}\int_{[0,1]^{\lvert{\mathfrak{u}}\rvert}}\left\lvert\int_{[0,1]^{s-\lvert{\mathfrak{u}}\rvert}}\left(\prod_{j\in{\mathfrak{u}}}\frac{\partial^{\alpha_{j}}}{\partial y_{j}^{\alpha_{j}}}\right)f({\boldsymbol{y}})\mathrm{d}{\boldsymbol{y}}_{-{\mathfrak{u}}}\right\rvert^{2}\mathrm{d}{\boldsymbol{y}}_{\mathfrak{u}},$
where $f:[0,1]^{d}\to{\mathbb{R}}$, ${\boldsymbol{y}}_{\mathfrak{u}}=({\boldsymbol{y}}_{j})_{j\in{\mathfrak{u}}}$, ${\boldsymbol{y}}_{-{\mathfrak{u}}}:=(y_{j})_{j\in{\mathcal{D}}\setminus{\mathfrak{u}}}$, and $\lvert{\mathfrak{u}}\rvert$ is the cardinality of ${\mathfrak{u}}$. The space $H$ is a special case of the weighted Korobov space which has real smoothness parameter $\alpha$ characterizing the rate of decay of Fourier coefficients, see e.g. (Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004).
The kernel (4) is the sum over $2^{d}$ terms and thus becomes impractical to compute for large $d$. A simplified space assumes $\gamma_{\mathfrak{u}}$ are product weights which take the form $\gamma_{\mathfrak{u}}=\prod_{j\in{\mathfrak{u}}}\gamma_{j}$ for some $\{\gamma_{j}\}_{j=1}^{d}$. The kernel with product weights takes the simplified form
(5) $K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)\right).$
Figure 3 shows one dimensional shift-invariant kernels with unit weight and varying smoothness. See (Kaarnioja et al., 2023) for a review of lattice-based kernel interpolation in such weighted spaces along with the development of serendipitous weights, which are a special form of high-performance product weights.
@@
so the Gram matrix ${\mathsf{K}}$ is circulant. Let $R_{m}(i)$ flip the first $m$ bits of $0\leq i<2^{m}$ in base $b=2$ so that if $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}2^{t}$ then $R(i)=\sum_{t=0}^{m-1}{\mathsf{i}}_{m-t-1}2^{t}$. The first step in the FFT with decimation-in-time is to reorder the inputs $\{y_{i}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$. The last step in computing the IFFT is to reorder the outputs $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{i}\}_{i=0}^{2^{m}-1}$. Therefore, one should generate lattice points in natural order and skip the first step of the FFT and last step of the inverse FFT.
The Fourier matrix is $\overline{{\mathsf{F}}^{(m)}}=\{W_{m}^{ij}\}_{i,j=0}^{2^{m}-1}$ where $W_{m}=\exp(-2\pi\sqrt{-1}/2^{m})$. Let $\overline{{\mathsf{T}}^{(m)}}=\{W_{m}^{iR_{m}(j)}\}_{i,j=0}^{2^{m}-1}$ so that
$\overline{{\mathsf{F}}^{(m)}\{y_{i}\}_{j=0}^{2^{m}-1}}=\overline{{\mathsf{T}}^{(m)}\{y_{R(j)}\}_{j=0}^{2^{m}-1}}.$
For $\{{\boldsymbol{x}}_{i}\}_{i=0}^{2^{m}-1}$ a lattice in natural order, we have
${\mathsf{K}}=\frac{1}{n}{\mathsf{T}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{T}}^{(m)}}.$
Notice that $\overline{{\mathsf{T}}^{(m+1)}}=\left\{W_{m+1}^{iR_{m+1}(j)}\right\}_{i,j=0}^{2^{m+1}-1}$. For $0\leq j<2^{m}$ we have $R_{m+1}(j)=2R_{m}(j)$ so $W^{iR_{m+1}(j)}_{m+1}=W^{iR_{m}(j)}_{m}$. For $2^{m}\leq j<2^{m+1}$ we have $R_{m+1}(j)=2R_{m}(j-2^{m})+1$ so $W^{iR_{m+1}(j)}_{m+1}=W_{m}^{iR(j-2^{m})}W_{m+1}^{i}$. Moreover, for $0\leq i<2^{m}$ we have $W_{m+1}^{2^{m}+i}=-W_{m+1}^{i}$. Define ${\widetilde{\boldsymbol{w}}}^{(m)}=\{W_{m+1}^{i}\}_{i=0}^{2^{m}-1}$. Then
$\overline{{\mathsf{T}}^{(m+1)}}=\begin{pmatrix}\overline{{\mathsf{T}}^{(m)}}&\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\\ \overline{{\mathsf{T}}^{(m)}}&-\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\end{pmatrix}$
and ${\widetilde{\boldsymbol{w}}}^{(m+1)}$ is the $\alpha=2$ interlacing of ${\widetilde{\boldsymbol{w}}}^{(m)}$ and $W_{m+1}{\widetilde{\boldsymbol{w}}}^{(m)}$.
5.2. Digital nets, digitally-shift-invariant kernels, and the FWHT
@@
denote digital addition and subtraction respectively. Note that in base $b=2$ these operations are both equivalent to the exclusive or (XOR) between bits. For vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in[0,1)^{d}$, digital operations ${\boldsymbol{x}}\oplus{\boldsymbol{y}}$ and ${\boldsymbol{x}}\ominus{\boldsymbol{y}}$ act component-wise. A kernel is said to be digitally-shift-invariant when
$K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}({\boldsymbol{u}}\ominus{\boldsymbol{v}})$
@@
Suppose $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a digitally shifted digital net in natural order (possibly of higher order and/or with LMS). Following (Jagadeeswaran and Hickernell, 2022, Theorem 5.3.1,Theorem 5.3.2), for $b=2$, the Gram matrix ${\mathsf{K}}$ is nested block Toeplitz which implies the eigendecomposition
${\mathsf{K}}=\frac{1}{n}{\mathsf{H}}^{(m)}{\mathsf{\Lambda}}^{(m)}{\mathsf{H}}^{(m)}$
where ${\mathsf{H}}^{(m)}$ is the $2^{m}\times 2^{m}$ Hadamard matrix defined by ${\mathsf{H}}^{(0)}=(1)$ and the relationship
${\mathsf{H}}^{(m+1)}=\begin{pmatrix}{\mathsf{H}}^{(m)}&{\mathsf{H}}^{(m)}\\ {\mathsf{H}}^{(m)}&-{\mathsf{H}}^{(m)}\end{pmatrix}.$
Multiplying by ${\mathsf{H}}^{(m)}$ can be done at at ${\mathcal{O}}(n\log n)$ cost using the FWHT. Note that ${\mathsf{K}}$ is not necessarily nested block Toeplitz under NUS.
We now describe one dimensional digitally-shift-invariant kernels $K(u,v)={\widetilde{K}}(u\ominus v)$ of higher-order smoothness $\alpha$. Weighted sums over products these one dimensional kernels may be used to construct higher-dimensional kernels e.g. those with product weights
$K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\left(K(u_{j},v_{j})-1\right)\right).$
To begin, let us write $k\in{\mathbb{N}}_{0}$ as $k=\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}b^{a_{\ell}}$ where $a_{1}>\cdots>a_{\#k}\geq 0$ are the #$\#k$ indices of non-zero digits ${\mathsf{k}}_{a_{\ell}}\in\{1,\dots,b-1\}$ in the base $b$ expansion of $k$. Then the $k^{\mathrm{th}}$ Walsh coefficient is
@@
For $\alpha>1$, let the digitally-shift-invariant kernel
(6) $K(x,z)=\sum_{k\in{\mathbb{N}}_{0}}\frac{\mathrm{wal}_{k}(x\ominus z)}{b^{\mu_{\alpha}(k)}}={\widetilde{K}}_{\alpha}(x\ominus z)$
have corresponding RKHS ${\widetilde{H}}_{\alpha}$. The following theorem shows ${\widetilde{H}}_{\alpha}$ contains smooth functions. The case when $\alpha=1$ is treated separately in (Dick and Pillichshammer, 2005).
@@
${\widetilde{H}}_{\alpha}\supset H_{\alpha}$ where $H_{\alpha}$ is an RKHS with inner product
$\langle f,g\rangle_{H_{\alpha}}=\sum_{l=1}^{\alpha-1}\int_{0}^{1}f^{(l)}(x)\mathrm{d}x\int_{0}^{1}g^{(l)}(x)\mathrm{d}x+\int_{0}^{1}f^{(\alpha)}(x)g^{(\alpha)}(x)\mathrm{d}x.$
@@
See appendix A
The following theorem gives explicit forms of a few higher-order digitally-shift-invariant kernels in base $b=2$ which are plotted in fig. 3. These kernels are also useful in the computation of the worst-case error of QMC with higher order polynomial-lattice rules as in (Baldeaux et al., 2012). In fact, their paper details how to compute ${\widetilde{K}}_{\alpha}$ at ${\mathcal{O}}(\alpha\#x)$ cost where #$\#x$ is the number of non-zero digits in the base $b$ expansion of $x$
@@
Fix the base $b=2$. Let $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ and $t_{\nu}(x)=2^{-\nu\beta(x)}$ where for $x=0$ we set $\beta(x)=t_{\nu}(x)=0$. Then
${\widetilde{K}}_{2}(x)$ $=-\beta(x)x+\frac{5}{2}\left[1-t_{1}(x)\right],$
${\widetilde{K}}_{3}(x)$ $=\beta(x)x^{2}-5\left[1-t_{1}(x)\right]x+\frac{43}{18}\left[1-t_{2}(x)\right],$
${\widetilde{K}}_{4}(x)$ $=-\frac{2}{3}\beta(x)x^{3}+5\left[1-t_{1}(x)\right]x^{2}-\frac{43}{9}\left[1-t_{2}(x)\right]x$
@@
The $\alpha\in\{2,3\}$ forms are due to (Baldeaux et al., 2012). The $\alpha=4$ form is derived in appendix A. ∎
5.3. Fast kernel computations and efficient updates
Let us make the general assumptions satisfied by the formulations in the previous two subsections.
  1. (1)
    For $n=2^{m}$, the Gram matrix ${\mathsf{K}}^{(m)}\in{\mathbb{R}}^{n\times n}$ can be written as
    ${\mathsf{K}}^{m}={\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{V}}^{(m)}}$

### v2

@@ 6.1 SI kernels, rank-1 lattices, and the FFTBR/IFFTBR
$K_{\alpha}(x,x^{\prime})=\frac{(2\pi)^{2\alpha}}{(-1)^{\alpha+1}(2\alpha)!}B_{2\alpha}((x-x^{\prime})\mod 1)$
with smoothness parameters ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$, global scale $\gamma>0$, weights (lengthscales) $\{\eta_{\mathfrak{u}}\}_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}$, and $B_{\ell}$ denoting the Bernoulli polynomial of degree $\ell$. The corresponding RKHS $H$ is a weighted periodic unanchored Sobolev space of smoothness ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$ with norm
$\lVert f\rVert_{H}^{2}:=\sum_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}\frac{1}{(2\pi)^{2\lvert{\mathfrak{u}}\rvert}\eta_{\mathfrak{u}}}\int_{[0,1]^{\lvert{\mathfrak{u}}\rvert}}\left\lvert\int_{[0,1]^{s-\lvert{\mathfrak{u}}\rvert}}\left(\prod_{j\in{\mathfrak{u}}}\frac{\partial^{\alpha_{j}}}{\partial y_{j}^{\alpha_{j}}}\right)f({\boldsymbol{y}})\mathrm{d}{\boldsymbol{y}}_{-{\mathfrak{u}}}\right\rvert^{2}\mathrm{d}{\boldsymbol{y}}_{\mathfrak{u}},$
where $f:[0,1]^{d}\to{\mathbb{R}}$, ${\boldsymbol{y}}_{\mathfrak{u}}=({\boldsymbol{y}}_{j})_{j\in{\mathfrak{u}}}$, ${\boldsymbol{y}}_{-{\mathfrak{u}}}:=(y_{j})_{j\in\{1,\dots,d\}\setminus{\mathfrak{u}}}$, and $\lvert{\mathfrak{u}}\rvert$ is the cardinality of $\mathfrak{u}$}. The space $H$ is a special case of the weighted Korobov space which has real smoothness parameter $\alpha$ characterizing the rate of decay of Fourier coefficients [Kaarnioja et al., 2022, 2023, Cools et al., 2021, 2020, Sloan and Woźniakowski, 2001, Kuo et al., 2004].
The kernel (5) is the sum over $2^{d}$ terms and thus becomes impractical to compute for large $d$. A simplified space assumes $\eta_{\mathfrak{u}}$ are product weights which take the form $\eta_{\mathfrak{u}}=\prod_{j\in{\mathfrak{u}}}\eta_{j}$ for some $\{\eta_{j}\}_{j=1}^{d}$. The kernel with product weights takes the simplified form
$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\prod_{j=1}^{d}\left(1+\eta_{j}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})\right).$ (6)
See [Kaarnioja et al., 2023] for a review of lattice-based kernel interpolation in such weighted spaces along with the development of serendipitous weights, which are a special form of high-performance product weights.
@@
so the Gram matrix $\mathsf{K}$} is circulant. Let $R_{m}(i)$ flip the first $m$ bits of $0\leq i<2^{m}$ in base $b=2$ so that if $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}2^{t}$ then $R(i)=\sum_{t=0}^{m-1}{\mathsf{i}}_{m-t-1}2^{t}$. The first step in the FFT with decimation-in-time is to reorder the inputs $\{y_{i}\}_{i=0}^{2^{m}-1}$ into bit-reversed order $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$. The last step in computing the IFFT is to reorder the outputs $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$ into bit-reversed order $\{y_{i}\}_{i=0}^{2^{m}-1}$. Therefore, one should generate lattice points in radical inverse order and skip the first step of the FFT and last step of the inverse FFT. We call such algorithms the FFTBR and IFFTBR respectively.
The Fourier matrix is $\overline{{\mathsf{F}}_{m}}=\{W_{m}^{ij}\}_{i,j=0}^{2^{m}-1}$ where $W_{m}=\exp(-2\pi\sqrt{-1}/2^{m})$. Let $\overline{{\mathsf{T}}_{m}}=\{W_{m}^{iR_{m}(j)}\}_{i,j=0}^{2^{m}-1}$ so that
$\overline{{\mathsf{F}}_{m}\{y_{i}\}_{j=0}^{2^{m}-1}}=\overline{{\mathsf{T}}_{m}\{y_{R(j)}\}_{j=0}^{2^{m}-1}}.$
For $\{{\boldsymbol{x}}_{i}\}_{i=0}^{2^{m}-1}$ a lattice in radical inverse order, we have
${\mathsf{K}}=\frac{1}{n}{\mathsf{T}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\mathsf{T}}_{m}}.$
Notice that $\overline{{\mathsf{T}}_{m+1}}=\left\{W_{m+1}^{iR_{m+1}(j)}\right\}_{i,j=0}^{2^{m+1}-1}$. For $0\leq j<2^{m}$ we have $R_{m+1}(j)=2R_{m}(j)$ so $W^{iR_{m+1}(j)}_{m+1}=W^{iR_{m}(j)}_{m}$. For $2^{m}\leq j<2^{m+1}$ we have $R_{m+1}(j)=2R_{m}(j-2^{m})+1$ so $W^{iR_{m+1}(j)}_{m+1}=W_{m}^{iR(j-2^{m})}W_{m+1}^{i}$. Moreover, for $0\leq i<2^{m}$ we have $W_{m+1}^{2^{m}+i}=-W_{m+1}^{i}$. Define ${\widetilde{\boldsymbol{w}}}_{m}=\{W_{m+1}^{i}\}_{i=0}^{2^{m}-1}$. Then
$\overline{{\mathsf{T}}_{m+1}}=\begin{pmatrix}\overline{{\mathsf{T}}_{m}}&\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})\overline{{\mathsf{T}}_{m}}\\ \overline{{\mathsf{T}}_{m}}&-\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})\overline{{\mathsf{T}}_{m}}\end{pmatrix}$
and ${\widetilde{\boldsymbol{w}}}_{m+1}$ is the $\alpha=2$ interlacing of ${\widetilde{\boldsymbol{w}}}_{m}$ and $W_{m+1}{\widetilde{\boldsymbol{w}}}_{m}$.
6.2 DSI kernels, digital nets, and the FWHT
@@
denote digital addition and subtraction respectively. Note that in base $b=2$ these operations are both equivalent to the exclusive or (XOR) between bits. For vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in[0,1)^{d}$, digital operations ${\boldsymbol{x}}\oplus{\boldsymbol{y}}$ and ${\boldsymbol{x}}\ominus{\boldsymbol{y}}$ act component-wise. A kernel is said to be digitally-shift-invariant (DSI) when
$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})={\widetilde{K}}({\boldsymbol{x}}\ominus{\boldsymbol{x}}^{\prime})$
@@
Suppose $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a digitally-shifted digital net in radical inverse order (possibly of higher-order and/or with LMS). Following [Rathinavel and Hickernell, 2022, Theorem 5.3.1,Theorem 5.3.2], for $b=2$, the Gram matrix $\mathsf{K}$} is RSBT (recursive symmetric block Toeplitz) which implies the eigendecomposition
${\mathsf{K}}=\frac{1}{n}{\mathsf{H}}_{m}{\mathsf{\Lambda}}_{m}{\mathsf{H}}_{m}$
where ${\mathsf{H}}_{m}$ is the $2^{m}\times 2^{m}$ Hadamard matrix defined by ${\mathsf{H}}^{(0)}=(1)$ and the relationship
${\mathsf{H}}_{m+1}=\begin{pmatrix}{\mathsf{H}}_{m}&{\mathsf{H}}_{m}\\ {\mathsf{H}}_{m}&-{\mathsf{H}}_{m}\end{pmatrix}$.
Multiplying by ${\mathsf{H}}_{m}$ can be done at ${\mathcal{O}}(n\log n)$ cost using the FWHT. Note that $\mathsf{K}$} is not generally RSBT when $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ has a NUS.
We now describe one dimensional DSI kernels $K_{\alpha}(x,x^{\prime})$ of higher-order smoothness $\alpha\geq 2$. Weighted sums over products these one dimensional kernels may be used to construct higher-dimensional kernels, e.g., those with product weights
$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\prod_{j=1}^{d}\left(1+\eta_{j}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})\right).$
To begin, let us write $k\in{\mathbb{N}}_{0}$ as $k=\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}b^{a_{\ell}}$ where $a_{1}>\cdots>a_{\#k}\geq 0$ are the #k\#k indices of non-zero digits ${\mathsf{k}}_{a_{\ell}}\in\{1,\dots,b-1\}$ in the base $b$ expansion of $k$. Then the $k^{\mathrm{th}}$ Walsh coefficient is
@@
For $\alpha\geq 2$, let the DSI kernel
$K_{\alpha}(x,x^{\prime})=\sum_{k\in{\mathbb{N}}}\frac{\mathrm{wal}_{k}(x\ominus x^{\prime})}{b^{\mu_{\alpha}(k)}}$ (7)
have corresponding RKHS $H_{\alpha}$. The following theorem shows $H_{\alpha}$ contains smooth functions. The case when $\alpha=1$ is treated separately in [Dick and Pillichshammer, 2005].
@@
Let $H_{\alpha}^{\prime}$ denote the RKHS with inner product
$\langle f,g\rangle_{H_{\alpha}^{\prime}}=\sum_{l=1}^{\alpha-1}\int_{0}^{1}f^{(l)}(x)\mathrm{d}x\int_{0}^{1}g^{(l)}(x)\mathrm{d}x+\int_{0}^{1}f^{(\alpha)}(x)g^{(\alpha)}(x)\mathrm{d}x$
and $f\in H_{\alpha}^{\prime}$ satisfying $\int_{0}^{1}f(x)\mathrm{d}x=0$. Then $H_{\alpha}^{\prime}\subset H_{\alpha}$ where $H_{\alpha}$ is the RKHS with kernel defined in Equation 7.
@@
See Appendix B
The following theorem gives explicit forms of a few higher-order DSI kernels in base $b=2$. These kernels are also useful in the computation of the worst-case error of QMC with higher-order polynomial-lattice rules as in [Baldeaux et al., 2012]. In fact, their paper details how to compute $K_{\alpha}$ in (7) at ${\mathcal{O}}(\alpha\#x)$ cost where #x\#x is the number of non-zero digits in the base $b$ expansion of $x$.
@@
Fix the base $b=2$. Let $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ and for $\nu\in{\mathbb{N}}$ define $t_{\nu}(x)=2^{-\nu\beta(x)}$ where for $x=0$ we set $\beta(x)=t_{\nu}(x)=0$. Then (7) satisfies
$K_{2}(x)$ $=-1+-\beta(x)x+\frac{5}{2}\left[1-t_{1}(x)\right],$
$K_{3}(x)$ $=-1+\beta(x)x^{2}-5\left[1-t_{1}(x)\right]x+\frac{43}{18}\left[1-t_{2}(x)\right],$
$K_{4}(x)$ $=-1+-\frac{2}{3}\beta(x)x^{3}+5\left[1-t_{1}(x)\right]x^{2}-\frac{43}{9}\left[1-t_{2}(x)\right]x$
@@
The $\alpha\in\{2,3\}$ forms are due to [Baldeaux et al., 2012]. The $\alpha=4$ form is derived in Appendix B. ∎
6.3 Fast kernel computations and efficient eigenvalue updates
The formulations in the previous two subsections have the following properties.
  1. 1.
    For $n=2^{m}$, the Gram matrix ${\mathsf{K}}_{m}\in{\mathbb{R}}^{n\times n}$ can be written as
    ${\mathsf{K}}_{m}={\mathsf{V}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\mathsf{V}}_{m}}$